# AAppendix References

### v3

@@ 3.2 Issuing Heat Alerts as a Constrained MDP
The heat alerts MDP components are summarized here. See also Table S1 of Appendix A. Each MDP ${\mathcal{M}}_{k}$ corresponds to a geographic area $k\in{\mathcal{K}}$, and episodes are indexed as $j\in{\mathcal{J}}$. In this work, geographic areas are U.S. counties, and each episode spans the warm season / summer from May 1st to September 30th of a specific year ($H=152$ days).
@@ 4.1 Data Sources
Here we provide an overview of the data used in the RL environment; Appendix B contains more details.
@@ 4.3 The Reward Function: A Hierarchical Model of Hospitalizations
To estimate the functions $\lambda_{k}$ and $\tau_{k}$ under limited data and low health signal, we incorporate a data-driven random effects prior (using the county-level features) and inject domain knowledge on the signs of certain coefficients. We train this model using variational inference, using the Python package Pyro (Harpole et al. 2019). The details of this approach, as well as model diagnostics, are in Appendix C. Note that the fitted coefficients in the alert effectiveness component indicate the presence of alert fatigue. To utilize this rewards model within the BROACH environment, we draw from the variational posterior distribution at the start of each episode, and then use those coefficients to calculate $r^{(k,j)}_{t}$ given $\big{(}s^{(k,j)}_{t},a^{(k,j)}_{t}\big{)}$ for $t=0,...,H-1$.
@@ 5.2 Experimental Setup
We investigate using four common RL algorithms: Deep Q-learning (dqn), Quantile Regression Deep Q-learning (qrdqn), Trust Region Policy Optimization (trpo), and Advantage Actor-Critic (a2c)—additional information on these algorithms is in Appendix D.2. We use standard implementations of these methods available in the Stable-Baselines3 Python library (Raffin et al. 2021). To ground our analysis and discussion, note that dqn and qrdqn are off-policy methods which learn deterministic policies by directly solving for the optimal value function using the Bellman optimality condition in equation 2; exploration is induced only during training (e.g. epsilon-greedy). Whereas, trpo and a2c are on-policy methods which learn stochastic policies by direct optimization of the expected return in equation 1, using refinements of the policy gradient theorem (Sutton et al. 1999); exploration is inherent.
@@
The counties for which nws issued few heat alerts had high-variance estimates of heat alert effectiveness in the rewards model. Therefore, for the RL experiments, we selected a subset of 30 counties in which at least 75 heat alerts were issued during 2006-2016; spread across the five major climate regions, prioritizing those with higher variance in estimated alert effectiveness across days. See map in Figure 2 and Appendix D.1 for more details.
@@
The main metric we use for policy comparison is the average cumulative reward per episode (“average return”). To estimate this metric fairly, we hold three years of data (2007, 2011, and 2015) out of RL training, referred to as the evaluation years [^6] Appendix D.4 shows the results of a sensitivity analysis where the models are trained on 2006-2013 and evaluated on 2014-2016.. For the purposes of RL hyperparameter tuning (details in Appendix D.3) and selection of the optimal QHI threshold for each county, we consider the regionally augmented weather trajectories and associated heat alert budgets from the evaluation years excluding the county of interest to be the validation data set. The final evaluation results (reported in tables and figures) are obtained by drawing 1,000 sets of coefficients from the Bayesian rewards model posterior and calculating the return under each policy with weather and budgets only from the county of interest during the evaluation years. These validation/tuning and evaluation procedures are described algorithmically in Appendix E. After calculating the average returns for the 30 counties under each competing policy, we compare each policy’s returns with the associated returns under the nws policy using a Wilcoxon-Mann-Whitney test. This nonparametric test allows us to compare the distribution of differences in returns, accounting for the fact that the counties are highly heterogeneous, so the differences between the policies are not normally distributed.
@@ 5.3 Results and Discussion
Figures 3, S5, and S6 illustrate how the different policies look in practice, respectively in scenarios where using RL is beneficial, where the alert budget is small, and where using RL is not beneficial compared to the alternatives.
@@
Several counterfactual policies perform worse than nws (as indicated by negative median differences in returns): random, basic.nws, and the RL models without the QHI restriction. By contrast, the RL models with the QHI restriction perform significantly better than nws, as indicated by their significant positive median difference in returns. A rough estimate if a2c.qhi were implemented as-is across all counties in the U.S. (using the Medicare population from 2011, the midpoint of our study period) is that we could see a reduction of 222 NOHR hospitalizations per year (approximate 95% CI = (-491, 1131); details of this CI calculation are in Appendix F). Discussion of this absolute benefit to public health is in Appendix G.
@@
Interpreting policies from the trpo and a2c models deterministically fails in the heat alerts setting under our rewards model. Additional discussion of this result is in Appendix H.2. This finding raises the question: Are stochastic policies palatable from a domain perspective? On the one hand, it is less immediately satisfying that some aspect of heat alerts issuance would be left to chance. On the other hand, in the context of human-in-the-loop decision making—which would likely be the case for public organizations such as NWS (Stuart et al. 2022)—an algorithm reporting probabilities is more informative than reporting only a binary action. In any case, if in the future a heat alerts RL model was running continuously online, it would likely need to utilize exploration (incorporating some randomness into its actions) to update itself over time.
@@
We note that the performance and behavior of dqn.qhi is practically identical to aa.qhi—not just in its numerical summary, but also in its policy behavior. qrdqn.qhi performs a bit better than dqn.qhi, but not as well as the stochastic policies. Between those, a2c.qhi performs better than trpo.qhi, though their policies tend to display similar characteristics. Therefore, we focus on a2c.qhi as the best RL policy in the post-hoc analysis. The non-RL counterfactual policies that perform well are that of always issuing an alert on days above an optimized QHI threshold (aa.qhi) and of issuing alerts on the $b$ hottest days of the summer (topk). We note that the ordering of the Wilcoxon-Mann-Whitney statistics for topk, aa.qhi, and a2c.qhi is opposite the ordering of these policies’ median difference in returns (compared to nws). This is due to the longer tails of the a2c.qhi differences, as illustrated in Figure 4. For interpreting these results, recall that topk is an oracle policy. However, aa.qhi is an implementable alternative to a2c.qhi, so we must consider it more seriously.
@@ 5.4 Post-hoc Contrastive Explanation
Figure 4 illustrates the varying performance of topk, aa.qhi, and a2c.qhi relative to the nws policy, across climate regions and counties. Note that large heterogeneity in the health effects of heat alert intervention across counties is consistent with previous findings (Wu et al. 2023). A striking visual pattern is the increasing vertical spread from left to right: topk performs the least heterogeneously across counties, followed by aa.qhi, and finally a2c.qhi. Similarly, the humid climate regions display more heterogeneity across counties than the others, across all the policies. This is partially due to those regions being larger and thus overrepresented among our 30 counties (see Table S3), but that does not fully explain the discrepancy.
Figure 4: Different policies’ average return relative to nws across different climate regions, over the evaluation years. Outlier points are labeled with their state abbreviation. Note that topk is an oracle policy.
@@ 6 Conclusion and Future Work Directions
Appendix
Appendix A Heat Alerts MDP Terminology
@@ B.3 Descriptive Statistics of Heat, Alerts, and Hospitalizations
Appendix C Details of the Bayesian Rewards Model
@@ D.4 Sensitivity Analysis Using 2014-2016 as the Evaluation Years
Appendix E Summary of the Workflow Using BROACH
@@
Appendix F Approximate Confidence Interval Calculation
@@
Appendix G Discussion of Absolute Health Benefits
@@
Appendix H Extended Results from the RL Analysis
@@ H.3 Additional Figures from the Post-hoc Contrastive Analysis
Figure S7 provides histograms of different policies’ temporal characteristics, and figures S8, S9, and S10 illustrate the CART analysis.

### v4

@@ 3.2 Issuing Heat Alerts as a Constrained MDP
The heat alerts MDP components are summarized here. See also Table S2 of Appendix A    Heat Alerts MDP Terminology. Each MDP ${\mathcal{M}}_{k}$ corresponds to a geographic area $k\in{\mathcal{K}}$, and episodes are indexed as $j\in{\mathcal{J}}$. In this work, geographic areas are U.S. counties, and each episode spans the warm season / summer from May 1st to September 30th of a specific year ($H=152$ days).
@@ 4.1 Data Sources
Here we provide an overview of the data used in the RL environment; Appendix B    Details of the Data Set contains more details.
@@ 4.3 The Reward Function: A Hierarchical Model of Hospitalizations
To estimate the functions $\lambda_{k}$ and $\tau_{k}$ under limited data and low health signal, we incorporate a data-driven random effects prior (using the county-level features) and inject domain knowledge on the signs of certain coefficients. We train this model using variational inference, using the Python package Pyro (Harpole et al. 2019). The details of this approach, as well as model diagnostics, are in Appendix C    Details of the Bayesian Rewards Model. Note that the fitted coefficients in the alert effectiveness component indicate the presence of alert fatigue. To utilize this rewards model within the BROACH environment, we draw from the variational posterior distribution at the start of each episode, and then use those coefficients to calculate $r^{(k,j)}_{t}$ given $\big{(}s^{(k,j)}_{t},a^{(k,j)}_{t}\big{)}$ for $t=0,...,H-1$.
@@ 5.2 Experimental Setup
We investigate using four common RL algorithms: Deep Q-learning (dqn), Quantile Regression Deep Q-learning (qrdqn), Trust Region Policy Optimization (trpo), and Advantage Actor-Critic (a2c)—additional information on these algorithms is in Appendix D.2    RL Algorithm Descriptions. We use standard implementations of these methods available in the Stable-Baselines3 Python library (Raffin et al. 2021). To ground our analysis and discussion, note that dqn and qrdqn are off-policy methods which learn deterministic policies by directly solving for the optimal value function using the Bellman optimality condition in equation 2; exploration is induced only during training (e.g. epsilon-greedy). Whereas, trpo and a2c are on-policy methods which learn stochastic policies by direct optimization of the expected return in equation 1, using refinements of the policy gradient theorem (Sutton et al. 1999); exploration is inherent.
@@
The counties for which nws issued few heat alerts had high-variance estimates of heat alert effectiveness in the rewards model. Therefore, for the RL experiments, we selected a subset of 30 counties in which at least 75 heat alerts were issued during 2006-2016; spread across the five major climate regions, prioritizing those with higher variance in estimated alert effectiveness across days. See map in Figure 2 and Appendix D.1    Selection of 30 Counties for more details.
@@
The main metric we use for policy comparison is the average cumulative reward per episode (“average return”). To estimate this metric fairly, we hold three years of data (2007, 2011, and 2015) out of RL training, referred to as the evaluation years [^6] Appendix D.4    Sensitivity Analysis Using 2014-2016 as the Evaluation Years shows the results of a sensitivity analysis where the models are trained on 2006-2013 and evaluated on 2014-2016.. For the purposes of RL hyperparameter tuning (details in Appendix D.3    Hyperparameter Tuning) and selection of the optimal QHI threshold for each county, we consider the regionally augmented weather trajectories and associated heat alert budgets from the evaluation years excluding the county of interest to be the validation data set. The final evaluation results (reported in tables and figures) are obtained by drawing 1,000 sets of coefficients from the Bayesian rewards model posterior and calculating the return under each policy with weather and budgets only from the county of interest during the evaluation years. These validation/tuning and evaluation procedures are described algorithmically in Appendix E    Summary of the Workflow Using BROACH. After calculating the average returns for the 30 counties under each competing policy, we compare each policy’s returns with the associated returns under the nws policy using a Wilcoxon-Mann-Whitney test. This nonparametric test allows us to compare the distribution of differences in returns, accounting for the fact that the counties are highly heterogeneous, so the differences between the policies are not normally distributed.
@@ 5.3 Results and Discussion
Figures 3, S8, and S9 illustrate how the different policies look in practice, respectively in scenarios where using RL is beneficial, where the alert budget is small, and where using RL is not beneficial compared to the alternatives.
@@
Several counterfactual policies perform worse than nws (as indicated by negative median differences in returns): random, basic.nws, and the RL models without the QHI restriction. By contrast, the RL models with the QHI restriction perform significantly better than nws, as indicated by their significant positive median difference in returns. A rough estimate if a2c.qhi were implemented as-is across all counties in the U.S. (using the Medicare population from 2011, the midpoint of our study period) is that we could see a reduction of 222 NOHR hospitalizations per year (approximate 95% CI = (-491, 1131); details of this CI calculation are in Appendix F    Approximate Confidence Interval Calculation). Discussion of this absolute benefit to public health is in Appendix G    Discussion of Absolute Health Benefits.
@@
Interpreting policies from the trpo and a2c models deterministically fails in the heat alerts setting under our rewards model. Additional discussion of this result is in Appendix H.2    Additional Discussion of Stochastic vs Deterministic Policies. This finding raises the question: Are stochastic policies palatable from a domain perspective? On the one hand, it is less immediately satisfying that some aspect of heat alerts issuance would be left to chance. On the other hand, in the context of human-in-the-loop decision making—which would likely be the case for public organizations such as NWS (Stuart et al. 2022)—an algorithm reporting probabilities is more informative than reporting only a binary action. In any case, if in the future a heat alerts RL model was running continuously online, it would likely need to utilize exploration (incorporating some randomness into its actions) to update itself over time.
@@
We note that the performance and behavior of dqn.qhi is practically identical to aa.qhi—not just in its numerical summary, but also in its policy behavior. qrdqn.qhi performs a bit better than dqn.qhi, but not as well as the stochastic policies. Between those, a2c.qhi performs better than trpo.qhi, though their policies tend to display similar characteristics. Therefore, we focus on a2c.qhi as the best RL policy in the post-hoc analysis. The non-RL counterfactual policies that perform well are that of always issuing an alert on days above an optimized QHI threshold (aa.qhi) and of issuing alerts on the $b$ hottest days of the summer (topk). We note that the ordering of the Wilcoxon-Mann-Whitney statistics for topk, aa.qhi, and a2c.qhi is opposite the ordering of these policies’ median difference in returns (compared to nws). This is due to the longer tails of the a2c.qhi differences, as illustrated in Figure S11. For interpreting these results, recall that topk is an oracle policy. However, aa.qhi is an implementable alternative to a2c.qhi, so we must consider it more seriously.
@@ 5.4 Post-hoc Contrastive Explanation
Figure S11 illustrates the varying performance of topk, aa.qhi, and a2c.qhi relative to the nws policy, across climate regions and counties. Note that large heterogeneity in the health effects of heat alert intervention across counties is consistent with previous findings (Wu et al. 2023). A striking visual pattern is the increasing vertical spread from left to right: topk performs the least heterogeneously across counties, followed by aa.qhi, and finally a2c.qhi. Similarly, the humid climate regions display more heterogeneity across counties than the others, across all the policies. This is partially due to those regions being larger and thus overrepresented among our 30 counties (see Table S4), but that does not fully explain the discrepancy.
@@ 6 Conclusion and Future Work Directions
Acknowledgements
This work was supported by an NSF Graduate Research Fellowship (EMC), NIH award K01ES032458 (RCN), NIH award R01-ES029950 (MT, GAW), NIH award 1R01MD016054-01A1 (FD), NIH award 5R01ES030616-04 (MT), NIH award 3RF1AG074372-01A1S1 (MT), NIH award P30ES000002 (MT), Alfred P. Sloan Foundation grant G-2020-13946 (FD), the Fernholz Foundation (MT), and Wellcome Trust grant 216033-Z-19-Z (GAW). We thank the National Studies on Air Pollution and Health research group for their support of this project. The computation in this paper was performed on the FASSE and FASRC Cannon clusters supported by the FAS Division of Science Research Computing at Harvard University. Lastly, we thank Kate Weinberger for facilitating access to the heat alerts data set, and the STAT 234 teaching staff at Harvard University for their feedback during early stages of this project.
Disclosures: GAW currently serves as a consultant for the Health Effects Institute (Boston, MA) and recently served as a consultant for Google, LLC (Mountain View, CA).
A    Heat Alerts MDP Terminology
@@ B.3    Descriptive Statistics of Heat, Alerts, and Hospitalizations
C    Details of the Bayesian Rewards Model
@@ D.4    Sensitivity Analysis Using 2014-2016 as the Evaluation Years
E    Summary of the Workflow Using BROACH
@@
F    Approximate Confidence Interval Calculation
@@
G    Discussion of Absolute Health Benefits
@@
H    Extended Results from the RL Analysis
@@ H.3    Additional Figures from the Post-hoc Contrastive Analysis
Figure S10 provides histograms of different policies’ temporal characteristics, figure S11 compares different policies’ performance (compared to NWS) across climate regions and counties, and figures S12, S13, and S14 illustrate the CART analysis.
@@
Figure S11: Different policies’ average return relative to nws across different climate regions, over the evaluation years. Outlier points are labeled with their state abbreviation. Note that topk is an oracle policy.